Calculate the potential difference of a battery connected to a capacitor consisting of two parallel plates with an area of 1.75 cm2 and a separation distance of 4.33 mm. Suppose that the charge present on the two plates is equal to 2.67 picocoulombs.

1
Expert's answer
2022-03-15T10:34:44-0400

The potential difference

"V=\frac{q}{C}=\frac{qd}{\epsilon_0A}"

"V=\frac{2.67*10^{-12}*4.33*10^{-3}}{8.85*10^{-12}*1.75*10^{-4}}=7.46\:\rm V"